# Learning Multiplex Embeddings on Text-rich Networks with One Text Encoder

### v1

@@ 4.1 Metern: Learning Multiplex Embeddings with One Text Encoder
We represent the relations by introducing relation prior tokens $P_{r}=\{p^{(1)}_{r},...,p^{(m)}_{r}\}$ with embeddings $Z_{r}=\{{\bm{z}}^{(1)}_{r},...,{\bm{z}}^{(m)}_{r}\}$ for each relation $r$. Each ${\bm{z}}^{(t)}_{r}\in Z_{r}$ is the $k$-dimensional embedding for the relation prior token $p^{(t)}_{r}\in P_{r}$. To generate relation-conditioned embeddings with one encoder, we provide the encoder with both the node information and the target relation, i.e., $d_{i}$ and $P_{r}$, as input. The representation ${\bm{h}}_{v_{i}|r}$ is obtained by
${\bm{h}}_{v_{i}|r}=\text{Enc}(\{P_{r};d_{i}\}),$ (2)
@@
$P_{r}(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}|r},{\bm{h}}_{v_{j}|r})=\text{Sim}(\text{Enc}(\{P_{r};d_{i}\}),\text{Enc}(\{P_{r};d_{j}\})).$ (3)
@@
$\max_{\Theta}\mathcal{O}=\sum_{r\in\mathcal{R}}\sum_{e_{ij}\in\mathcal{E}^{r}}\text{log}P_{r}(e_{ij}|v_{i},v_{j};\Theta),$ (4)
Here, the conditional probability $P_{r}(e_{ij}|v_{i},v_{j};\Theta)$ is calculated as follows Oord et al. (2018):
$P_{r}(e_{ij}|v_{i},v_{j};\Theta)=\frac{\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{j}|r})}{\sum_{v_{u}\in\mathcal{V}}\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{u}|r})},$ (5)
To make the calculation efficient, we leverage the negative sampling technique Jin et al. (2020); Mikolov et al. (2013) to simplify the objective and obtain our loss below.
$\min_{\Theta}\mathcal{L}=\sum_{r\in\mathcal{R}}w_{r}\sum_{e_{ij}\in\mathcal{E}^{r}}-\log\frac{\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{j}|r})}{\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{j}|r})+\sum_{v_{u}^{\prime}}\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{u}^{\prime}|r})}=\sum_{r\in\mathcal{R}}w_{r}\mathcal{L}_{r}.$ (6)
In the equation above, $v_{u}^{\prime}$ stands for a random negative sample. In our implementation, we use “in-batch negative samples” Karpukhin et al. (2020) to reduce the encoding cost. Note that we add relation weights $w_{r}$ to control the relative learning speed of different relations (analysis can be found in Section 5.8).
4.2 Inference with Metern
Node embeddings of high quality should be able to generalize to various downstream tasks within the given network scenario Hamilton et al. (2017). In this section, we propose to conduct direct inference and “learn-to-select-source-relations” inference with the learned multiplex embeddings.
@@ 4.2 Inference with Metern
${\bm{h}}_{\text{target}}={\bm{h}}_{v|r_{\text{target}}}=\text{Enc}(\{P_{r_{\text{target}}};d\}).$ (7)
@@
We propose to learn to select source relations with some downstream task training samples when the semantically closest source relation is not clear (e.g., paper classification and year prediction in academic networks; item classification and price prediction in e-commerce networks). We introduce a set of learnable query embeddings for the target task $Q_{\text{target}}=\{{\bm{q}}^{(1)}_{\text{target}},{\bm{q}}^{(2)}_{\text{target}},...,{\bm{q}}^{(s)}_{\text{target}}\}$ to learn to select the source relation embeddings from $Z_{\mathcal{R}}=\{Z_{r}\}_{r\in\mathcal{R}}$ via attention-based mix-up, as follows:
${\bm{z}}^{(t)}_{\text{target}}=\sum_{{\bm{z}}_{i}\in Z_{\mathcal{R}}}\alpha^{(t)}_{i}\cdot{\bm{z}}_{i},\quad\quad\alpha^{(t)}_{i}=\text{softmax}_{{\bm{z}}_{i}\in Z_{\mathcal{R}}}({\bm{z}}_{i}\cdot{\bm{q}}^{(t)}_{\text{target}})$ (8)
$Z_{\text{target}}=\{{\bm{z}}^{(1)}_{\text{target}},{\bm{z}}^{(2)}_{\text{target}},...,{\bm{z}}^{(s)}_{\text{target}}\},\quad\quad P_{\text{target}}=\{p^{(1)}_{\text{target}},p^{(2)}_{\text{target}},...,p^{(s)}_{\text{target}}\},$ (9)
@@
$Z_{\mathcal{R}}=\{Z_{r}\}_{r\in\mathcal{R}}$ is the source relation embedding pool (with ||×$|\mathcal{R}|\times m$ embeddings), which is the union of all the source relation embeddings for $r\in\mathcal{R}$. We use $P_{\text{target}}$ and $Z_{\text{target}}$ to represent the downstream task prior tokens and their embeddings respectively (${\bm{z}}^{(t)}_{\text{target}}$ is the embedding for $p^{(t)}_{\text{target}}$). Note that only the parameters in $Q_{\text{target}}$ are trainable to learn to select source relations.
@@ 4.3 Discussions
In our experiment, we use the Transformer encoder Vaswani et al. (2017) as the $\text{Enc}(\cdot)$ function. Time Complexity: Given a node $v_{i}$ associated with a document $d_{i}$ containing $p$ tokens, the time complexity of our multiplex representation learning model with $m$ embeddings for each relation is $\mathcal{O}((p+m)^{2})$, which is on par with the $\mathcal{O}(p^{2})$ complexity of the single relation representation learning model since $m\ll p$. Memory Complexity: Given a network with $|\mathcal{R}|$ types of relations and $T$ parameters in the $\text{Enc}(\cdot)$, the parameter complexity of our multiplex representation learning model is $\mathcal{O}(T+|\mathcal{R}|mk)$, which is nearly the same with the complexity of $\mathcal{O}(T)$ of the single relation representation learning model since $|\mathcal{R}|mk\ll T$. The empirical time efficiency study and memory efficiency study are shown in Section 5.6.
Difference from Existing Works Lester et al. (2021); Liu et al. (2021); Qin and Eisner (2021).
These works propose to add learnable “prompt tokens” onto the text sequences, before feeding them into the Transformer architecture, sharing a similar design with our relation prior tokens. However, there are three inherent differences between these works and our work: 1) Different focuses: Existing works Lester et al. (2021); Liu et al. (2021); Qin and Eisner (2021) focus more on efficient language model tuning (i.e., training the prompt embedding only), while we focus on joint relation learning (i.e., using the relation embeddings to capture relation specific information and the shared encoder to capture cross-relation shared knowledge). 2) Different designs: The prompt tokens used in NLP problems Lester et al. (2021); Liu et al. (2021); Qin and Eisner (2021) generally correspond to natural language descriptions of the tasks and labels. In our case, we use relation prior tokens to encode the abstract relations between documents that may not be directly described by natural language tokens. As a result, the representation of an unseen relation may need to be learned as a mixture of source relation representations instead of simply being represented by natural language prompts. 3) Different applications: Existing works focus on NLU Liu et al. (2021); Qin and Eisner (2021), NLG Lester et al. (2021), and knowledge probe Qin and Eisner (2021) tasks, while our work focuses on multiplex representation learning.
@@ 5 Experiments
In this section, we first introduce the five datasets. Then, we demonstrate the high quality of the learned multiplex embeddings by Metern. After that, we show the effectiveness of Metern on downstream tasks with direct source relation inference and “learn-to-select-source-relation” inference, respectively. Finally, we conduct efficiency analysis, multiplex embedding visualization, and the study of relation weight $w_{r}$. More experiments on relation token embedding initialization and relation token embedding visualization can be found in Appendix A.6 and A.8, respectively.
5.1 Datasets
We run experiments on both academic networks from the Microsoft Academic Graph (MAG) Sinha et al. (2015); Zhang et al. (2023) and e-commerce networks from Amazon He and McAuley (2016). In academic networks, nodes correspond to papers and there are five types of relations among papers: “cited-by” (cb), “same-author” (sa), “same-venue” (sv), “co-reference” (cr), and “co-cited-by” (ccb); while in e-commerce networks, nodes are items and there are four types of relations between items: “co-purchased” (cop), “co-viewed” (cov), “bought-together” (bt), and “co-brand” (cob). Since both MAG and Amazon have multiple domains, we select two domains from MAG and three domains from Amazon. In total, we have five datasets in the experiments (i.e., MAG-Geology, MAG-Mathematics, Amazon-Clothes, Amazon-Home, and Amazon-Sports). The datasets’ statistics can be found in Table 7.
5.2 Baselines
We compare Metern with three kinds of baselines, large-scale corpora finetuned text embedders, multiplex graph neural networks, and multi-relation learning language models. The first category includes SPECTER Cohan et al. (2020), SciNCL Ostendorff et al. (2022), Sentence-Transformer Reimers and Gurevych (2019), and the OpenAI-ada-002 embedder Brown et al. (2020). SPECTER Cohan et al. (2020) is a text embedder finetuned from SciBERT Beltagy et al. (2019) with academic domain contrastive learning. SciNCL Ostendorff et al. (2022) further improves SPECTER by introducing controlled nearest neighbor sampling. For Sentence-Transformer Reimers and Gurevych (2019), we use the recent best checkpoint [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 which finetunes MPNet Song et al. (2020) on above 1 billion sentence pairs from 32 domains. OpenAI-ada-002333https://openai.com/blog/new-and-improved-embedding-model is the recent large language model-based text embedder proposed by OpenAI. The second category includes DMGI Park et al. (2020) and HDMI Jing et al. (2021). DMGI Park et al. (2020) is a graph convolutional network-based multiplex node embedding encoder. HDMI Jing et al. (2021) further improves DMGI by proposing an alternative semantic attention-based fusion module. For multiplex GNN methods, we use bag-of-words embeddings as the initial node feature vectors. The third category includes the vanilla finetuned language model (Vanilla FT) and MTDNN Liu et al. (2019a). For Vanilla FT, we finetuned BERT Devlin et al. (2019) with all the relation pairs, without distinguishing among them. In this way, the model will only output one single embedding for each node. MTDNN Liu et al. (2019a) is a multi-task learning model with a shared language model backbone for different tasks and task-specific modules connected to the backbone. As a result, it can output multiplex representations for each node corresponding to different tasks (relations). Note that Vanilla FT, MTDNN, and Metern are all initialized by the same bert-base-uncased checkpoint.
Table 1: Multiplex representation learning experiments on academic networks: Geology and Mathematics. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively.
@@ 5.2 Baselines
Table 2: Multiplex representation learning experiments on e-commerce networks: Clothes, Home, and Sports. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively.
@@ 5.3 Multiplex Representation Learning
We test the quality of the generated embeddings of different models on multiplex relation prediction tasks on the network. Given a query node/text (e.g., a paper or an item), a target relation (e.g., cb, sa, cop, or cov), and a candidate node/text list (e.g., papers or items), we aim to predict which key node in the candidate list should be linked to the given query node under the target relation. We use PREC@1 as the metric. More detailed information on experimental settings can be found in Appendix A.5.1
The results on academic networks and e-commerce networks are shown in Table 1 and Table 2, respectively. From the result, we can find that: 1) Metern performs significantly and consistently better than all the baseline methods on all types of relation prediction on all datasets (except cop prediction on Home and Sports). 2) In academic networks, multiplex representation learning methods (MTDNN and Metern) can always outperform the single representation learning method (Vanilla FT); while in e-commerce networks, Vanilla FT performs on par with MTDNN and Metern. This is because the relations in e-commerce networks are more semantically close to each other, while the relations in academic networks are less close to or even conflict with each other (see Figure 2 and Figure 7). 3) Although SPECTER Cohan et al. (2020), SciNCL Ostendorff et al. (2022), Sentence-Transformer Reimers and Gurevych (2019), and OpenAI-ada-002 Brown et al. (2020) are finetuned on large-scale corpora, they perform badly compared with Metern since they only output one embedding which cannot capture the diverse relation semantics. 4) Multiplex GNN methods (DMGI and HDMI) can utilize neighbor information from the network to output multiplex embeddings but fail to capture the diverse and dynamic text semantics associated with nodes.

### v2

@@ 4.1 Metag: Learning Multiplex Embeddings with One Language Model Encoder
We represent the relations by introducing relation prior tokens $P_{r}=\{p^{(1)}_{r},...,p^{(m)}_{r}\}$ with embeddings $Z_{r}=\{{\bm{z}}^{(1)}_{r},...,{\bm{z}}^{(m)}_{r}\}$ for each relation $r$. Each ${\bm{z}}^{(t)}_{r}\in Z_{r}$ is a $k$-dimensional embedding vector for the relation prior token $p^{(t)}_{r}\in P_{r}$. To generate relation-conditioned embeddings with one encoder, we provide the encoder with both the node information and the target relation, i.e., $d_{i}$ and $P_{r}$, as input. The representation ${\bm{h}}_{v_{i}|r}$ is obtained by
$\small{\bm{h}}_{v_{i}|r}=\text{Enc}(\{P_{r};d_{i}\}),$ (3)
@@
$\small P_{r}(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}|r},{\bm{h}}_{v_{j}|r})$ (4)
$=\text{Sim}(\text{Enc}(\{P_{r};d_{i}\}),\text{Enc}(\{P_{r};d_{j}\})).$ (5)
@@
$\small\max_{\Theta}\mathcal{O}=\sum_{r\in\mathcal{R}}\sum_{e_{ij}\in\mathcal{E}^{r}}\text{log}P_{r}(e_{ij}|v_{i},v_{j};\Theta),$ (6)
Here, the conditional probability $P_{r}(e_{ij}|v_{i},v_{j};\Theta)$ is calculated as follows (Oord et al., 2018) ($s_{v_{i},v_{j}}={\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{j}|r}$):
$\small P_{r}(e_{ij}|v_{i},v_{j};\Theta)=\frac{\exp(s_{v_{i},v_{j}})}{\sum_{v_{u}\in\mathcal{V}}\exp(s_{v_{i},v_{u}})}.$ (7)
To make the calculation efficient, we leverage the negative sampling technique (Jin et al., 2020; Mikolov et al., 2013) to simplify the objective and obtain our loss below.
$\scriptsize\mathcal{L}=\sum_{r\in\mathcal{R}}w_{r}\sum_{e_{ij}\in\mathcal{E}^{r}}-\log\frac{\exp(s_{v_{i},v_{j}})}{\exp(s_{v_{i},v_{j}})+\sum_{v_{u}^{\prime}}\exp(s_{v_{i},v_{u}^{\prime}}))}.$ (8)
In the equation above, $v_{u}^{\prime}$ stands for a random negative sample. In our implementation, we use “in-batch negative samples” (Karpukhin et al., 2020) to reduce the encoding cost. Note that we add relation weights $w_{r}$ to control the relative learning speed of different relations (analysis can be found in Section 5.8).
4.2 Inference with Metag
Node embeddings of high quality should be able to generalize to various downstream tasks within the given graph scenario (Hamilton et al., 2017). In this section, we propose to conduct direct inference and “learn-to-select-source-relations” inference with the learned multiplex embeddings.
@@ 4.2 Inference with Metag
$\small{\bm{h}}_{\text{target}}={\bm{h}}_{v|r_{\text{target}}}=\text{Enc}(\{P_{r_{\text{target}}};d\}).$ (9)
@@
We propose to learn to select source relations with some downstream task training samples when the semantically closest source relation is not clear (e.g., paper classification and year prediction in academic graphs; item classification and price prediction in e-commerce graphs). We introduce a set of learnable query embeddings for the target task $Q_{\text{target}}=\{{\bm{q}}^{(1)}_{\text{target}},{\bm{q}}^{(2)}_{\text{target}},...,{\bm{q}}^{(s)}_{\text{target}}\}$ to learn to select the source relation embeddings from $Z_{\mathcal{R}}=\{Z_{r}\}_{r\in\mathcal{R}}$ via attention-based mix-up, as follows:
${\bm{z}}^{(t)}_{\text{target}}=\sum_{{\bm{z}}_{i}\in Z_{\mathcal{R}}}\alpha^{(t)}_{i}\cdot{\bm{z}}_{i},\,\alpha^{(t)}_{i}=\text{softmax}_{{\bm{z}}_{i}\in Z_{\mathcal{R}}}({\bm{z}}_{i}\cdot{\bm{q}}^{(t)}_{\text{target}})$ (10)
$Z_{\text{target}}=\{{\bm{z}}^{(1)}_{\text{target}},...,{\bm{z}}^{(s)}_{\text{target}}\},\,P_{\text{target}}=\{p^{(1)}_{\text{target}},...,p^{(s)}_{\text{target}}\}$ (11)
@@
$Z_{\mathcal{R}}=\{Z_{r}\}_{r\in\mathcal{R}}$ is the source relation embedding pool (with ||×$|\mathcal{R}|\times m$ embeddings), which is the union of all the source relation embeddings for $r\in\mathcal{R}$. We use $P_{\text{target}}$ and $Z_{\text{target}}$ to represent the downstream task prior tokens and their embeddings, respectively (i.e., ${\bm{z}}^{(t)}_{\text{target}}$ is the embedding for $p^{(t)}_{\text{target}}$). Note that only the parameters in $Q_{\text{target}}$ are trainable to learn to select source relations.
@@ 4.3 Discussions
In our experiment, we use the Transformer encoder (Vaswani et al., 2017) as the $\text{Enc}(\cdot)$ function. Time Complexity: Given a node $v_{i}$ associated with a document $d_{i}$ containing $p$ tokens, the time complexity of our multiplex representation learning model with $m$ embeddings for each relation is $\mathcal{O}((p+m)^{2})$, which is on par with the $\mathcal{O}(p^{2})$ complexity of the single relation representation learning model since $m\ll p$. Memory Complexity: Given a graph with $|\mathcal{R}|$ types of relations and $T$ parameters in the $\text{Enc}(\cdot)$, the parameter complexity of our multiplex representation learning model is $\mathcal{O}(T+|\mathcal{R}|mk)$, which is nearly the same as the $\mathcal{O}(T)$ complexity of the single relation representation learning model since $|\mathcal{R}|mk\ll T$. The empirical time efficiency study and memory efficiency study are shown in Section 5.6.
Difference from Existing Works (Lester et al., 2021; Liu et al., 2021; Qin & Eisner, 2021).
These works propose to add learnable “prompt tokens” to the text sequences before feeding them into the Transformer architecture, sharing a similar design with our relation prior tokens. However, there are three inherent differences between these works and our work: (1) Different focuses: Existing works (Lester et al., 2021; Liu et al., 2021; Qin & Eisner, 2021) focus more on efficient language model tuning (i.e., training the prompt embedding only), while we focus on joint representation learning (i.e., using the relation embeddings to capture relation-specific information and the shared encoder to capture knowledge shared across relations). (2) Different designs: The prompt tokens used in NLP problems (Lester et al., 2021; Liu et al., 2021; Qin & Eisner, 2021) generally correspond to natural language descriptions of the tasks and labels. In our case, we use relation prior tokens to encode the abstract relations between documents that may not be directly described by natural language tokens. As a result, the representation of an unseen relation may need to be learned as a mixture of source relation representations instead of simply being represented by natural language prompts. (3) Different applications: Existing works focus on NLU (Liu et al., 2021; Qin & Eisner, 2021), NLG (Lester et al., 2021), and knowledge probing (Qin & Eisner, 2021) tasks, while our work focuses on multiplex representation learning.
@@ 5 Experiments
Table 1: Multiplex representation learning experiments on academic graphs: Geology and Mathematics. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation, respectively.
@@
Table 2: Multiplex representation learning experiments on e-commerce graphs: Clothes, Home, and Sports. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation, respectively.
@@
In this section, we first introduce the five datasets. Then, we demonstrate the high quality of the learned multiplex embeddings by Metag. After that, we show the effectiveness of Metag on downstream tasks with direct source relation inference and “learn-to-select-source-relation” inference, respectively. Finally, we conduct efficiency analysis, multiplex embedding visualization, and the study of the relation weight $w_{r}$. More experiments on relation token embedding initialization and relation token embedding visualization can be found in Appendices A.6 and A.8, respectively.
5.1 Datasets
We run experiments on both academic graphs from the Microsoft Academic Graph (MAG) (Sinha et al., 2015; Zhang et al., 2023) and e-commerce graphs from Amazon (He & McAuley, 2016). In academic graphs, nodes correspond to papers and there are five types of relations among papers: “cited-by” (cb), “same-author” (sa), “same-venue” (sv), “co-reference” (cr), and “co-cited-by” (ccb); while in e-commerce graphs, nodes are items and there are four types of relations between items: “co-purchased” (cop), “co-viewed” (cov), “bought-together” (bt), and “co-brand” (cob). Since both MAG and Amazon have multiple domains, we select two domains from MAG and three domains from Amazon. In total, we have five datasets in the experiments (i.e., MAG-Geology, MAG-Mathematics, Amazon-Clothes, Amazon-Home, and Amazon-Sports). The datasets’ statistics can be found in Appendix Table 7.
5.2 Baselines
We compare Metag with three kinds of baselines, large-scale corpora finetuned text embedders, multiplex graph neural networks, and multi-relation learning language models. The first category includes SPECTER (Cohan et al., 2020), SciNCL (Ostendorff et al., 2022), Sentence-Transformer (Reimers & Gurevych, 2019), and the OpenAI-ada-002 embedder (Brown et al., 2020). SPECTER (Cohan et al., 2020) is a text embedder finetuned from SciBERT (Beltagy et al., 2019) with citation-guided contrastive learning. SciNCL (Ostendorff et al., 2022) further improves SPECTER by introducing controlled nearest neighbor sampling. For Sentence-Transformer (Reimers & Gurevych, 2019), we use the recent best checkpoint [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 which finetunes MPNet (Song et al., 2020) with over 1 billion sentence pairs from 32 domains. OpenAI-ada-002333https://openai.com/blog/new-and-improved-embedding-model is the recent large language model-based text embedder proposed by OpenAI. The second category includes DMGI (Park et al., 2020) and HDMI (Jing et al., 2021). DMGI (Park et al., 2020) is a graph convolutional network-based multiplex node embedding encoder. HDMI (Jing et al., 2021) further improves DMGI by proposing an alternative semantic attention-based fusion module. For multiplex GNN methods, we use bag-of-words embeddings as the initial node feature vectors. The third category includes the vanilla finetuned language model (Vanilla FT) and MTDNN (Liu et al., 2019a). For Vanilla FT, we finetuned BERT (Devlin et al., 2019) with all the relation pairs, without distinguishing among them. In this way, the model will only output one single embedding for each node. MTDNN (Liu et al., 2019a) is a multi-task learning model with a shared language model backbone for different tasks and task-specific modules connected to the backbone. As a result, it can output multiplex representations for each node corresponding to different tasks (relations). Note that Vanilla FT, MTDNN, and Metag are all initialized by the same bert-base-uncased checkpoint.
@@ 5.3 Multiplex Representation Learning
We test the quality of the generated embeddings of different models on multiplex relation semantic matching tasks on the graphs. Given a query node/text (e.g., a paper or an item), a target relation (e.g., cb, sa, cop, or cov), and a candidate node/text list (e.g., papers or items), we aim to predict which key node in the candidate list should be linked to the given query node under the target relation semantics. We use PREC@1 as the metric. More detailed information on experimental settings can be found in Appendix A.5.1.
The results on academic graphs and e-commerce graphs are shown in Table 1 and Table 2, respectively. From the result, we can find that: 1) Metag performs significantly and consistently better than all the baseline methods on all types of relation prediction on all datasets (except cop prediction on Home and Sports). 2) In academic graphs, multiplex representation learning methods (MTDNN and Metag) can always outperform the single representation learning method (Vanilla FT); while in e-commerce graphs, Vanilla FT performs on par with MTDNN and Metag. This is because the relations in e-commerce graphs are semantically closer to each other, while the relations in academic graphs are more dissimilar or even conflict with each other (see Figure 1 and Figure 7). 3) Although SPECTER (Cohan et al., 2020), SciNCL (Ostendorff et al., 2022), Sentence-Transformer (Reimers & Gurevych, 2019), and OpenAI-ada-002 (Brown et al., 2020) are finetuned on large-scale corpora, they perform badly compared with Metag since they only output one embedding which cannot capture the diverse relation semantics. 4) Multiplex GNN methods (DMGI and HDMI) can utilize neighbor information from the graph to output multiplex embeddings but fail to capture the diverse and contextualized text semantics associated with nodes. More experiments can be found in Appendix A.9, A.10, and A.11.